# Large deviations for the smallest eigenvalue of a deformed GOE with an outlier

### v1

@@ 2.6. Solving the functional equation in the region $\omega_{\nu,t}(\lambda)\leq\Lambda$
Since $\Lambda_{1}\neq\alpha$, it follows that $\Lambda_{1}$ equals $\omega_{\nu,t}^{*}(\lambda)$.
We can now give the proof of Lemma 2.8. Let us prove the upper bound (2.29). Let $Y(\gamma)\in\Delta^{p+1}$ be given by (2.31) with $\gamma$ as in (2.27). By (2.32), we know that $K_{\nu,t}(\Lambda,\lambda,Y(\gamma))=I_{\nu,t}^{\Lambda}(\lambda).$ Using (2.25), and taking $Y=Y(\gamma)$ in the infimum yields
@@
which gives $F_{\nu,t}^{+}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda)$ using (2.29). Since $F_{\nu,t}^{-}(\Lambda,\lambda)\geq F_{\nu,t}^{+}(\Lambda,\lambda)$ we get that $F_{\nu,t}^{-}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda)$.
@@ 2.7. Solving the functional inequalities
In the next lemma we prove that the equality (2.30) in fact holds for any $\Lambda\leq\ell_{\nu}$ and $\lambda\leq\ell_{\nu,t}$, hence completing the proof of the weak large deviation principle (2.3).
@@
Let $\Lambda\leq\ell_{\nu}$. Let $\Phi(\Lambda,Y)$ be as in (2.12). Let $I_{\nu,t}^{\Lambda}$ be as in (2.28). Let $\lambda<\ell_{\nu,t}$. Let $\omega_{\nu,t}^{*}(\lambda)$ be the unique solution $\gamma$ in $[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu})$ of the equation $H_{\nu,t}(\gamma)=\lambda$. Define
$\mathcal{B}(\Lambda,\lambda):=\inf_{Y\in\Delta^{p+1}}\Bigr{\{}K_{\nu,t}(\Lambda,\lambda,Y)+I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,Y)}\Bigr{\}}.$
Then, the function $K_{\nu,t}(\Lambda,\lambda,\cdot)+I_{\nu,t}^{\Phi(\Lambda,\cdot)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,\cdot)}$ admits a unique minimizer $Y^{*}\in\Delta^{p+1}$. Furthermore,
@@
$\mathcal{B}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda).$
@@
Since $I_{\nu,t}^{\Phi(\Lambda,Y)}\geq 0$, we deduce from the first step of the proof of Lemma 2.9 that $G$ is bounded from below (by $K_{\nu,t}$). We claim that it is lower semicontinuous in $Y\in\Delta^{p+1}$ when $\Delta^{p+1}$ is equipped with the weak topology. We already checked in the previous proof that $J_{\nu,t}^{\Lambda}(\lambda,.)$ is lower semicontinuous . Moreover $Y\mapsto\Phi(\Lambda,Y)$ is continuous by Lemma 2.4. Furthermore, when $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\lambda)$, we have
@@
$-\lambda\eta_{k}+\eta_{k}^{2}-R_{1}\eta_{k}-\frac{t}{y_{k}}-C_{0}\frac{1}{\eta_{k}-\Lambda_{1}}=R_{2}$
@@
with $R_{1}=\Lambda y_{0}+\int\eta_{\nu}\mathrm{d}\mu$, ${C_{0}=\partial_{\Lambda}I_{\nu,t}^{\Phi(\Lambda,Y)}(.,\lambda)|_{\Phi(\Lambda,Y)}}$ and $\Lambda_{1}=\Phi(\Lambda,Y)$. It follows that for each $k=1,\ldots,p$,
$y_{k}=\frac{t}{\eta_{k}^{2}-(\lambda+R_{1})\eta_{\nu}-R_{2}-\frac{C_{0}}{\eta_{k}-\Lambda_{1}}}=\frac{t(\eta_{k}-\Lambda_{1})}{(\eta_{k}^{2}-(\lambda+R_{1})\eta_{k}-R_{2})(\eta_{\nu}-\Lambda_{1})-C_{0}}.$
@@
$\frac{y_{0}}{\Lambda-\Lambda_{1}}+\sum_{k=1}^{p}\frac{t}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=0.$ (2.49)
@@
$y_{0}+t\sum_{k=1}^{p}\frac{\eta_{k}-\Lambda_{1}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}\mathrm{d}x=1,$
@@
$t\sum_{k=1}^{p}\frac{1}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}+t(\Lambda-\Lambda_{1})\sum_{k=1}^{p}\frac{1}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=1-y_{0}.$
@@
$\sum_{k=1}^{p}\frac{t}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=1.$ (2.51)
@@
$\Lambda y_{0}+t\sum_{k=1}^{p}\frac{(\eta_{k}-\Lambda_{1})\eta_{k}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=R_{1}=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1},$
@@
$\Lambda y_{0}+t\sum_{k=1}^{p}\frac{\eta_{k}-\Lambda_{1}}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}+\Lambda(1-y_{0})=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1}.$
@@
$y_{k}=\frac{(\eta_{k}-\Lambda_{1})}{(\eta_{k}-\Lambda)(\eta_{k}-\omega_{\nu,t}(\lambda))(\eta_{k}-\omega_{\nu,t}^{*}(\lambda))},\quad\text{for each $1\leq k\leq p$},$ (2.54)
@@
Let $(y_{0},Y)$ be of the form (2.54). Let us compute $G(y_{0},Y)$. Define $X:=\Lambda y_{0}+\sum_{k=1}^{p}\eta_{k}y_{k}$. Recall that $X=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1}$. We have
$\sum_{k=1}^{p}\eta_{k}^{2}y_{k}=t-(\Lambda_{1}-\Lambda)tG_{\nu}(\Lambda)+(\gamma_{1}+\gamma_{2})X-\Lambda(\gamma_{1}+\gamma_{2})y_{0}-\gamma_{1}\gamma_{2}(1-y_{0}).$
Therefore
$\Lambda^{2}y_{0}+\sum_{k=1}^{p}\eta_{k}^{2}y_{k}=(\Lambda-\gamma_{1})(\Lambda-\gamma_{2})y_{0}-\gamma_{1}\gamma_{2}-(\Lambda_{1}-\Lambda)tG_{\nu}(\Lambda)+(\gamma_{1}+\gamma_{2})X.$
Inserting (2.53), we can therefore rewrite the above display as
$\Lambda^{2}y_{0}+\sum_{k=1}^{p}\eta_{k}^{2}y_{k}=-\gamma_{1}\gamma_{2}+(\gamma_{1}+\gamma_{2})X+(\Lambda-\Lambda_{1})(\Lambda-\lambda).$
It follows that
$\begin{split}G(y_{0},Y)&=\frac{1}{2}-\frac{1}{2}\ln t+\frac{1}{4t}(-2\lambda-X+2(\gamma_{1}+\gamma_{2}))X-\frac{\gamma_{1}\gamma_{2}}{2t}+\frac{1}{2t}(\Lambda-\Lambda_{1})(\Lambda-\lambda)\\ &-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2})-S_{\nu}(\Lambda_{1}))+\frac{1}{2}(S_{\nu}(\gamma_{1})-S_{\nu}(\Lambda_{1}))+\frac{1}{4t}(\Lambda_{1}^{2}-2\lambda\Lambda_{1}-\gamma_{1}^{2}+2\lambda\gamma_{1})\\ &=-\frac{1}{2}(S_{\nu}(\gamma_{2})+S_{\nu}(\Lambda))+\frac{Y}{4t}+\frac{1}{2}-\frac{1}{2}\ln t,\end{split}$
@@
$\begin{split}Y&:=(-2\lambda-X+2(\gamma_{1}+\gamma_{2}))X-2\gamma_{1}\gamma_{2}+2(\Lambda^{2}+\Lambda_{1}\lambda-\Lambda\lambda-\Lambda\lambda)+\Lambda_{1}^{2}-2\lambda\Lambda_{1}-\gamma_{1}^{2}+2\lambda\gamma_{1}\\ &=-\Lambda^{2}+\lambda^{2}+\gamma_{1}^{2}+\gamma_{2}^{2}-2\lambda\gamma_{1}-2\lambda\gamma_{2}+2\gamma_{1}\gamma_{2}+2\Lambda\Lambda_{1}-\Lambda_{1}^{2}-2\gamma_{1}\gamma_{2}\\ &+2(\Lambda^{2}+\Lambda_{1}\lambda-\Lambda\lambda-\Lambda\lambda)+\Lambda_{1}^{2}-2\lambda\Lambda_{1}-\gamma_{1}^{2}+2\lambda\gamma_{1}\\ &=\Lambda^{2}+\lambda^{2}+\gamma_{2}^{2}-2\gamma_{2}\lambda-2\lambda\Lambda.\end{split}$
@@
By Lemma 2.8, we only need to consider the case where $\Lambda<\omega_{\nu,t}(\lambda)$ and $\lambda<\ell_{\nu,t}$. Let $\lambda<\ell_{\nu,t}$ and $\Lambda<\omega_{\nu,t}(\lambda)$. Recall
@@
Recall that for every $\Lambda,Y$, $\Phi(\Lambda,Y)\geq\Lambda$. Consequently for each $k\geq 1$ and all $Y_{1},\ldots,Y_{k}\in\Delta^{p+1}$,
@@
  • Or $y_{0}\leq\delta/\gamma$ and denote by $\tilde{Y}$ the element of $\Delta^{p+1}$ such that $\tilde{y}_{0}=0$ and $\tilde{y}_{i}=y_{i}(1-y_{0})^{-1}$ for $i\in\{1,\ldots,p\}$. The continuity of $K_{\nu,t}(\Lambda,\lambda,.)$ implies that for $\delta$ small enough, uniformly on $Y,\Lambda,\lambda$,
@@
    Notice that the right-hand side does not depend on $\Lambda$ and moreover by (2.55), $I_{\nu,t}^{\omega^{*}_{\nu,t}(\lambda))}(\lambda)>0$. We hereafter choose $\delta$ small enough such that $\kappa(\lambda):=\min_{Y\in\Delta^{p}}\{K_{\nu,t}(\Lambda,\lambda,(0,Y))\}+O(\delta)$ is positive and independent of $\Lambda$.
Returning to (2.60), and recalling that $(Y_{1}^{\varepsilon},\ldots,Y_{n}^{\varepsilon})$ is a minimizer, and that $F_{\nu,t}^{+}$ is lower semicontinuous in the variable $\Lambda$ while $K_{\nu,t}$ and $\omega_{\nu,t}$ are continuous, we obtain the following: if again $(Y_{1},\cdots,Y_{n})$ denotes a limiting point of the optimizing sequence $(Y_{1}^{\varepsilon},\ldots,Y_{n}^{\varepsilon})$,
@@
which is impossible for $n$ large enough since $F_{\nu,t}^{+}(\Lambda,\lambda)$ is finite. Therefore, we deduce that there exists $n$ large enough, such that for each $k\geq n$, $\Phi_{k}(\Lambda,Y_{1},\ldots,Y_{k})>\omega_{\nu,t}(\lambda)$. We can argue as in the first step to conclude that

### v2

@@ 2.6. Solving the functional equation in the complement of the region $\Lambda<\omega_{\nu,t}(\lambda)$
Since $\Lambda_{1}\neq\alpha$, it follows that $\Lambda_{1}$ equals $\omega_{\nu,t}^{*}(\lambda)$. The proof of (2) is immediate. ∎
We can now give the proof of Lemma 2.8.
Proof of Lemma 2.8.
Let us prove the upper bound (2.37). Let $Y(\gamma)\in\Delta^{p+1}$ be given by (2.39) with $\gamma$ as in (2.35). By (2.40), we know that $K_{\nu,t}(\Lambda,\lambda,Y(\gamma))=I_{\nu,t}^{\Lambda}(\lambda).$ Using (2.33), and taking $Y=Y(\gamma)$ in the infimum yields
@@
which gives $F_{\nu,t}^{+}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda)$ using (2.37). Since $F_{\nu,t}^{-}(\Lambda,\lambda)\geq F_{\nu,t}^{+}(\Lambda,\lambda)$ we get that $F_{\nu,t}^{-}(\Lambda,\lambda)=I_{\nu,t}^{\Lambda}(\lambda)$. ∎
@@ 2.7. Solving the functional inequalities
In the next lemma, we prove that the equality (2.38) in fact holds for any $\Lambda\leq\ell_{\nu}$ and $\lambda\leq\ell_{\nu,t}$. In view of Lemma 2.8, it only remains to prove it for $\Lambda<\omega_{\nu,t}(\lambda)$, i.e in case $C$ of Figure 2. This will complete the proof of the weak large deviation principle (2.3).
@@
Let $\Lambda\leq\ell_{\nu}$. Let $\Phi(\Lambda,Y)$ be as in (2.15), $\rho_{\nu,t}^{\Lambda}$ be as in (1.13) and $\omega_{\nu,t}^{*}$ be as (1.15). Let $I_{\nu,t}^{\Lambda}$ be as in (2.36). Let $\lambda<\ell_{\nu,t}$. Let $\omega_{\nu,t}^{*}(\lambda)$ be the unique solution $\gamma$ in $[\omega_{\nu,t}(\ell_{\nu,t}),\ell_{\nu})$ of the equation $H_{\nu,t}(\gamma)=\lambda$. Then, the function $K_{\nu,t}(\Lambda,\lambda,\cdot)+I_{\nu,t}^{\Phi(\Lambda,\cdot)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,\cdot)}$ admits a unique minimizer $Y^{*}\in\Delta^{p+1}$. Furthermore,
@@
$I_{\nu,t}^{\Lambda}(\lambda)=\mathcal{B}(\Lambda,\lambda):=\inf_{Y\in\Delta^{p+1}}\Bigr{\{}K_{\nu,t}(\Lambda,\lambda,Y)+I_{\nu,t}^{\Phi(\Lambda,Y)}(\lambda)1_{\omega_{\nu,t}(\lambda)\geq\Phi(\Lambda,Y)}\Bigr{\}}.$
@@
Since $I_{\nu,t}^{\Phi(\Lambda,Y)}\geq 0$, we deduce from the first step of the proof of Lemma 2.9 that $G$ is bounded from below (by $K_{\nu,t}$). We claim that it is lower semicontinuous in $Y\in\Delta^{p+1}$ when $\Delta^{p+1}$ is equipped with the weak topology. We already checked in the previous proof that $J_{\nu,t}^{\Lambda}(\lambda,.)$ is lower semicontinuous . Moreover $Y\mapsto\Phi(\Lambda,Y)$ is continuous by Lemma 2.4. Furthermore, when $\Phi(\Lambda,Y)\leq\omega_{\nu,t}(\lambda)$, we have by Lemma 2.8
@@
$-\lambda\eta_{k}+\eta_{k}^{2}-R_{1}\eta_{k}-\frac{t\alpha_{k}}{y_{k}}-C_{0}\frac{1}{\eta_{k}-\Lambda_{1}}=R_{2}$
@@
with $R_{1}=\Lambda y_{0}+\int\eta_{\nu}\mathrm{d}\mu$, ${C_{0}=\partial_{\Lambda}I_{\nu,t}^{\Phi(\Lambda,Y)}(.,\lambda)|_{\Phi(\Lambda,Y)}}$ and $\Lambda_{1}=\Phi(\Lambda,Y)$. It follows that for every $k=1,\ldots,p$,
$y_{k}=\frac{t\alpha_{k}}{\eta_{k}^{2}-(\lambda+R_{1})\eta_{\nu}-R_{2}-\frac{C_{0}}{\eta_{k}-\Lambda_{1}}}=\frac{t\alpha_{k}(\eta_{k}-\Lambda_{1})}{(\eta_{k}^{2}-(\lambda+R_{1})\eta_{k}-R_{2})(\eta_{\nu}-\Lambda_{1})-C_{0}}.$
@@
$\frac{y_{0}}{\Lambda-\Lambda_{1}}+\sum_{k=1}^{p}\frac{t\alpha_{k}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=0.$ (2.59)
@@
$y_{0}+t\sum_{k=1}^{p}\alpha_{k}\frac{\eta_{k}-\Lambda_{1}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=1,$
@@
$t\sum_{k=1}^{p}\frac{\alpha_{k}}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}+t(\Lambda-\Lambda_{1})\sum_{k=1}^{p}\frac{\alpha_{k}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=1-y_{0}.$
@@
$\sum_{k=1}^{p}\frac{t\alpha_{k}}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=1.$ (2.61)
@@
$\Lambda y_{0}+t\sum_{k=1}^{p}\frac{\alpha_{k}(\eta_{k}-\Lambda_{1})\eta_{k}}{(\eta_{k}-\Lambda)(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}=R_{1}=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1},$
@@
$\Lambda y_{0}+t\sum_{k=1}^{p}\frac{\alpha_{k}(\eta_{k}-\Lambda_{1})}{(\eta_{k}-\gamma_{1})(\eta_{k}-\gamma_{2})}+\Lambda(1-y_{0})=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1}.$
@@
$y_{k}=\frac{\alpha_{k}(\eta_{k}-\Lambda_{1})}{(\eta_{k}-\Lambda)(\eta_{k}-\omega_{\nu,t}(\lambda))(\eta_{k}-\omega_{\nu,t}^{*}(\lambda))},\quad\text{{for every} $1\leq k\leq p$},$ (2.64)
@@
Let $(y_{0},Y)$ be of the form (2.64). Let us compute $G(y_{0},Y)$. Define $M(\gamma):=\Lambda y_{0}+\sum_{k=1}^{p}\eta_{k}y_{k}$. Recall that $M(\gamma)=\Lambda+\gamma_{1}+\gamma_{2}-\lambda-\Lambda_{1}$. We have
$\begin{split}\sigma(\gamma)&:=\Lambda^{2}y_{0}+\sum_{k=1}^{p}\eta_{k}^{2}y_{k}=\Lambda^{2}y_{0}+t-(\Lambda_{1}-\Lambda)tG_{\nu}(\Lambda)+(\gamma_{1}+\gamma_{2})X-\Lambda(\gamma_{1}+\gamma_{2})y_{0}-\gamma_{1}\gamma_{2}(1-y_{0})\\ &=(\Lambda-\gamma_{1})(\Lambda-\gamma_{2})y_{0}-\gamma_{1}\gamma_{2}-(\Lambda_{1}-\Lambda)tG_{\nu}(\Lambda)+(\gamma_{1}+\gamma_{2})M(\gamma)\\ &=\gamma_{1}\gamma_{2}+(\gamma_{1}+\gamma_{2})M(\gamma)+(\Lambda-\Lambda_{1})(\Lambda-\lambda).\end{split}$
where we inserted (2.63). It follows that
$\begin{split}G(y_{0},Y)&=\frac{1}{2}-\frac{1}{2}\ln t+\frac{1}{4t}(-2\lambda-M(\gamma)+2(\gamma_{1}+\gamma_{2}))M(\gamma)-\frac{\gamma_{1}\gamma_{2}}{2t}+\frac{1}{2t}(\Lambda-\Lambda_{1})(\Lambda-\lambda)\\ &-\frac{1}{2}(S_{\nu}(\Lambda)+S_{\nu}(\gamma_{1})+S_{\nu}(\gamma_{2})-S_{\nu}(\Lambda_{1}))+\frac{1}{2}(S_{\nu}(\gamma_{1})-S_{\nu}(\Lambda_{1}))+\frac{1}{4t}(\Lambda_{1}^{2}-2\lambda\Lambda_{1}-\gamma_{1}^{2}+2\lambda\gamma_{1})\\ &=-\frac{1}{2}(S_{\nu}(\gamma_{2})+S_{\nu}(\Lambda))+\frac{Z}{4t}+\frac{1}{2}-\frac{1}{2}\ln t,\end{split}$
@@
$\begin{split}Z&:=(-2\lambda-M(\gamma)+2(\gamma_{1}+\gamma_{2}))M(\gamma)-2\gamma_{1}\gamma_{2}+2(\Lambda^{2}+\Lambda_{1}\lambda-\Lambda\lambda-\Lambda\lambda)+\Lambda_{1}^{2}-2\lambda\Lambda_{1}-\gamma_{1}^{2}+2\lambda\gamma_{1}\\ &=\Lambda^{2}+\lambda^{2}+\gamma_{2}^{2}-2\gamma_{2}\lambda-2\lambda\Lambda.\end{split}$
@@
By Lemma 2.8, we only need to consider the case where $\Lambda<\omega_{\nu,t}(\lambda)$ and $\lambda<\ell_{\nu,t}$ which we assume in the sequel. Recall
@@
Recall that for every $\Lambda,Y$, $\Phi(\Lambda,Y)\geq\Lambda$. Consequently for every $k\geq 1$ and all $Y_{1},\ldots,Y_{k}\in\Delta^{p+1}$,
@@
  • Or $y_{0}\leq\delta/\gamma$. Then denote by $\tilde{Y}$ the element of $\Delta^{p+1}$ such that $\tilde{y}_{0}=0$ and $\tilde{y}_{i}=y_{i}(1-y_{0})^{-1}$ for $i\in\{1,\ldots,p\}$. The continuity of $K_{\nu,t}(\Lambda,\lambda,.)$ implies that for $\delta$ small enough, uniformly on $Y,\Lambda,\lambda$,
@@
    Notice that the right-hand side does not depend on $\Lambda$ and moreover by (2.65), $I_{\nu,t}^{\omega^{*}_{\nu,t}(\lambda))}(\lambda)>0$. We hereafter choose $\delta$ small enough such that
    $\kappa(\lambda):=\min_{Y\in\Delta^{p}}\{K_{\nu,t}(\Lambda,\lambda,(0,Y))\}+O(\delta)$
    is strictly positive and independent of $\Lambda$.
Returning to (2.70), and recalling that $(Y_{1}^{\varepsilon},\ldots,Y_{n}^{\varepsilon})$ is a minimizer, and that $F_{\nu,t}^{+}$ is lower semicontinuous in the variable $\Lambda$ while $K_{\nu,t}$ and $\omega_{\nu,t}$ are continuous, we obtain the following: if again $(Y_{1},\cdots,Y_{n})$ denotes a limiting point of the optimizing sequence $(Y_{1}^{\varepsilon},\ldots,Y_{n}^{\varepsilon})$, then
@@
which is impossible for $n$ large enough since $F_{\nu,t}^{+}(\Lambda,\lambda)$ is finite. Therefore, we deduce that there exists $n$ large enough, such that for every $k\geq n$, $\Phi_{k}(\Lambda,Y_{1},\ldots,Y_{k})>\omega_{\nu,t}(\lambda)$. We can argue as in the first case to conclude that